E-time and D-time for 20190505 UT

Summary

Tried TCS 2019A, but had problems. Switched back to 2018B, but had problems starting up.

AO OSCO at the beginning closed AO loop on both sides.

PEPSIPOL Active Optics on DX was finally solved with WFS_FlipX=True since wfs camera installed in spring for LUCI1 uses alternate amplifier.

Successful imaging of two geostationary satellites using LBC with NS hijacks.

Tests of LUCI acquisition and guiding with "fast" objects like Quasi.

Attempted SOUL ESM observations stopped by problems with SX Bayside stages. (IT 7682) Collected some DX ESM data.

Preparations

Test TCS Build 2019A

As mentioned this morning, we have switched TCS build from 2018B to 2019A at 00:45 UT. The first problem encountered was that MCSGUI would not launch. From a command line it prints "MCSGUI Version 3.5", but no GUI appears. The second problem encountered was that OSSGUI reports "MCSPU Error", and the error persists after we restart both MCSPU and the OSS subsystem. Thus at 01:05 UT we started switching back to 2018B. Instruments which don't auto-connect to TCS/IIF should now be reconnected. It was a great training exercise for Josh. We now return you to your normal D/E observing.......with TCS Build 2018B.

01:45 Can't launch ECSGUI in 2018B - Josh calls software person and they do alot of restarting on tcs3.

Details

AO OSCO with LUCI/LUCI (Josh & Doug)

02:55 UT - Josh is doing DX OSCO, while Doug is running SX SOUL (notes from each of them are elsewhere)

03:19 UT - Authorize LUCI2 only as SX SOUL WFS is having issues.

03:40 UT - Authorize as LUCI/LUCI to give SX SOUL another try. AO loops got closed on both sides.

04:00 UT - Begin switchover to PEPSIPOL/PEPSIPOL

PEPSIPOL Active Optics Alignment (John)

04:31 UT - ACTIVE preset to binary star HIP 57639 = HD 102616

Captured output from IDL's gcs_wfsc_display, /show_pupil, /show_center is in the two attached files from user lbto. Initial DX collimation was done with IDL.

Flipped sign of Z5,Z6, Z8, on DX AGW4 in PEPSIPOL_R_RM_sub11.cfg but why did this undo some of the last adjustments I did last night?

04:41 UT - restart GCSR and resend ACTIVE preset

Active optics for SX PEPSIPOL with AGw3 continues to be stable. Runaway problems on DX. Flipped sign of Z7, Z9 (all by changing sextScale_Zx config parameters)

04:47 UT - restart GCSR and resend ACTIVE preset (restarting GCS isn't needed for these changes, *sub11.cfg is read on the fly)

Flipped sign of Z5, z10. GCSR will converge sometimes, but signs are different as we change rotator angle.

Josh measures SX -7, -10 DEXTY and DX -6, -8 DETXY to offset star onto fiber, but on wrong component of binary star.

ACTIVE preset to single star so that things are more clear.

Z6 sign changes again ???????????????????

Summary: Probe offset need to put star on fiber varies from target to target, so seems to be some kind of transform problem. We don't really have time to sort that out, so we'll rely on Ilya sending offset telescope commands to center things up for now.

Summary: GCS DX active optics is only semi-stable - sign of Z6 and other Zernikes is flipping back and forth with rotator angle or something.
GCS and IDL say they are rotating by exactly the same angle. *IDL is flipping Z5, Z8, Z10.*
How can that be? Wait. Eureka.....IDL is flipping the X-zernikes, but GCSR is not. This is because of using the alternate amplifier in AGW4 wfs since last spring, and I had not applied the WFS_flipX parameter to PEPSIPOL_R.cfg.

05:34 UT - Applied WFS_flipX True in PEPSIPOL_R.cfg, and put several of my other sign flips back to normal. Now DX active optics seems to work correctly. Thanks to Ilya for doing 10 minutes of offset tests which gave me time to think about this solution.

05:46 UT Switching to LBC/LBC for the next project.

Geostationary Satellite Imaging with LBC (Olga & John)

06:05 Slew to copointing field for fainter satellite (B~14.6)

06:09 Running dohybrid

~06:20 John found the target GALAXY in JPL/Horizons and we are sending a NS preset now. Target saturated in 1-sec at V, dropped exptime by factor 2, but upon pause and resume, NS preset was lost. (Resuming from a paused state resends the preset?)

06:30 Trying again - now with 0.3-sec. b 063324 r 063343. The scanning, Bonn, shutter will mean that for these short exptimes one side of the field is exposed before the other. The astrometry would have to take this into account, but header information indicates the direction of the scan. This set was with BV and 0.3-sec.

06:42 Now trying a set with 1-sec exposures and Uz. Non-sidereal tracking using the JPL/Horizons ephemeris. Object is saturated in 1-sec at z-SLOAN.

06:46 Now trying with 1-sec and UY. At Y the target is not saturated, peak counts ~45000. Collected over 10 images on the blue and red sides in this configuration.

07:00 Slew to the brighter target ANIK. Use UY already and 0.1-sec exposure times. Also use the NS hijack since the ephemeris of this satellite is in JPL/Horizons.

07:08 Take a sidereal exposure of 0.1-sec at UY. b 07084? r 070840 . Nothing in these fields.

07:13:35 Send another sidereal preset, this time 1-sec and for coords at 07:15. I had turned off telescope & tracking, so we took data but did not execute a preset (we tracked sidereally at old coords, I think?)

07:20 Now with tel & tracking on, I sent preset to take a sidereal image. Shutter opened just after 07:20, though coords of target were for 07:20. First exposures are r 072041 and b 072048. There are 2 and 3 trails on the E side of the images and the rest are stars.

07:24 Now a non-sidereal preset, hijacked using NSIGUI. Exptimes 0.1-sec with U and Y filters. There are a few NS targets here, it seems. Stars are not very trailed in these short 0.1-sec exposures. Have ~15 0.1-sec exposures. Will take some longer exposures as well.

07:34 A new non-sidereal preset to obtain a series of 1-sec exposures through U and Y. Look OK - this target is not as bright as I thought it would be, however it is well detected, along with a number of stars.

LUCI/LUCI tests (Olga & Barry)

07:44 Josh is reconfiguring to LUCI1/LUCI2. We'll skip the LUCI/LBC and the GCS/WFS tests and go on to the non-linearity. The skies are pretty clear, although there may be some cirrus on the horizon.

08:10 L2 MOS error and I had to call Dave. He is working on it. Barry is here for the AO/ESM/SOUL work.

08:21 Dave resolved the MOS error, but meanwhile we determined that there was too much cirrus around for the non-linearity test.

08:30 Going to try the WCS/GCS test on quasi. Opening the LUCI observer interface from /home/lbto/igor/luci to the new RTD version that handles NS acquisitions. Also opening the GCS GUI Guide Control windows.

08:32 Set the hijack and send the LUCI script. After the preset was sent (though I think it could have been done before it was sent), I set the WFS exposure time to 4000 ms and clicked "set". Current WFS time is set to 4000 ms on left and was 4000 ms also on the right, although it then went up to 18000 sec. The star is not well collimated. Times keep going up and GCSGUI reads: "Blocked: try to shorten expsoure time". Clicking set and it says "unblocking as exposure time changed." But sometimes it seems first to read once more "Blocked..." before clearing. The autoscale of exposure time, however, happens after every image, it seems. The guide star had R=15.03, however. With a brighter guide star, perhaps the autoscaling would not have happened.

08:46 Lost left WFS image - spots are no longer there. guide star image just froze. But it revived at 08:46:58 - only guide cam, WFS image has no spots. Time to limit is 55 (according to GCS gui, but also from the script prep, I think there is room and we're not near the edge of the patrol field), so I don't know why the left WFS is showing no spots.

08:55 Aligning using the new RTD. The "Set" buttons under Target1, Target2 are small/invisible (this was an issue in the past and has re-emerged with this version). Otherwise, it seems to work fine. The offsets it calculates are what we would expect given the pixel scales and the X,Y positions of target and slit. We sent the offsets but the object is hard to see and we did not try hard to see it in the thru-slit image.

ESM tests (Doug & Barry)

09:11 Moving on to the ESM dither test. Barry is calling Doug and Josh will be resetting both SX and DX AdSecs for ESM.

09:32 Sending a LUCI2 ESM preset. AOref star R=10.76, LUCIGS R=13.79. Paused after the first exposure to try to adjust field stop position (since we had noticed that it was misaligned during the previous test), but could not see edges even at K. It is OK as is.

10:37 Authorizing LUCI1/LUCI2 for a SOUL/SX+ESM/DX test.

UT 10:39 - moving to zenith to set flats and reauthorize to binocular
UT 10:40 - at zenith authorize for LUCI-Bino
UT 10:42 - sending preset to M92 - AO SX - ESM -DX
UT 10:46 - waiting for IIF on SX - although collimation is done on both sides
- Preset is not completeing on AO - SX side

UT 10:52 - D. Miller killed SX preset
-Bayside stages are not moving on SX-AGW-1

UT 11:04 - D. Miller powering off the SX bayside stages - we will do
SL on SX and ESM on DX data collection for remainder of night - IT 7682.

-will start with R=9.4 AO Ref star for DX
UT 11:10 - sending preset for M92 SL-SX ESM-DX DIMM 0.85"
UT 11:11 - right side failed - resending
UT 11:18 - shell RIP'd
UT 11:25 - resending preset
UT 11:33 - DIMM 0.76" - SX WFS 0.6"
0007-0014 L1 0038-0045 L2 Ks-band
luci2.20190505.0038 - PSFs elongated (1st exposure)

0015-0023 L1 0046-0054 H-band DIMM 0.61"

UT 11:50 - L1 H-band images look almost blurry - L2 looks great

luci1.20190505.0015 - looks out of focus, but next one is better

UT 12:05 - stopping script - closing up will take flats

UT 12:19 - starting cals

L1

0024-0028 Hflats lights off

0029-0032 Hflats lights on

0033-0037 Ks lights off

0038-0042 Ks lights on

L2

0055-0059 Hflats lights off

0060-0064 Hflats lights on

0065-0069 Ks lights off

0070-0074 Ks lights on

UT 12:28 - finished with cals - putting LUCIs in safe config.


Planned: Engineering and D-time Tasks for (UT) May 5, 17, 22, and 23
  1. DONE PEPSI on-sky checks - IT 7545 - 1 hour (May 5 - first half of night)
    • conditions: seeing up to 1.5, clouds OK
    • Sort out DX active optics parameters for GCSR - See IT 7545
    • Work out PCS hotspot offsets to put start onto PEPSIPOL fiber
  2. DONE Non-Sidereal Tests - GCS/WFS tests with LUCI-1/LUCI-2 - 30 min (any)
    • Can be folded in with #10 below
  3. DONE LUCI-2 ESM Dithering Tests - 45 min (any)
    • RA 14hr
  4. LUCI-1/LUCI-2 Non-linearity on-sky tests - 30 min (May 5)
    • RA 17hr (IC 4665)
  5. LUCI/LBC mixed-mode test (any)
  6. ESM vs SL data collection - any amount of time - at least 30 min (May 5 or May 21/22)
    • DONE M92 RA - 17hr - H and K only
    • HCG 57 RA - 11hr
    • Hoag's Object RA - 15hr
  7. AGW-1 On-sky checkout (May 17 or May 22)
  8. LUCI-1/AGW-1 transform data collection (May 22)
  9. MODS On-sky checkout (May 22)
  10. AO on-sky checkout (May 22)
  11. DONE IT 7642 - NS acquisition - update to RTD - test the update during normal operations to make sure no unforseen side effects occur (all) - background (any)
    • There is a local version of the LUCI RTD that has a fix for this in the obs machines as user "lbto" under
      /home/lbto/igor/luci
    • and in the remote room machines as user "observer" under
      /home/observer/igor/luci
    • once you are in that directory simply run
      ./open_observer
      or more directly
      ./LUCI_Observer
      make sure you have the dot before the "/" so it picks the local one.
  12. DONE Continued GCS/PCS/TCS tests - background
    • activate TCS build 2019A, let it run (only TCS; MCSPU and OACserver are already running new version)
    • there might be problems with OSS (M2&M3) UMACs. If possible, please activate after handover and do some test presets, watching OSS for M2 and M3 control (and possible errors)
  13. D-time Science Programs
    • LBC test program 2hours (May 5 only) - requires clear sky and good seeing (faint targets)
    • ESM/LUCI-2 imaging of Triple Merger (K-band) ~45min-1hr (start of night May 5 only)
  14. Poor weather backup program - any amount of time - targets cover entire night
    • H-band library - K and M stars
    • H-band library - case example of QSO rest-frame 1.6micron CO-bandhead

Night Plan for UT May 05

  1. DONE On-Sky-CheckOut (OSCO) of both LUCI-1 SX/SOUL / LUCI-2 DX - Twilight
    • Need ~30min to check performance - this might extend slightly past twilight
  2. DONE PEPSI on-sky checkout - 2-hours - 8:30pm
  3. DONE LBC geostationary test - 1-2 hours - 10:30-11:30pm
  4. LUCI/LBC Mixed Mode (any field)
  5. LUCI-Bino NS GSC/WFS test (needs fast-moving NS target. Target used last time will be available until 2am) - 30 min -
    • DONE GCS/WFS test- fast mover/bright - check reduced exposure times, auto checks can be overridden - This REQUIRES using the NSIGUI (target: 469219 or 2016 HO3)
      • rm507-N:~/Service_Observing/E_IT7643WFS/
      • --- location of scripts (binoLUCI), README and OT program file.The scripts are based on UT time
      • Will need to manually change GCS GUI exposure times
      • open the GCS GUI, in the "Window" menu at the top left
      • select "Guide Control"
      • Near the bottom of the window that pops up next to "Nominal WFS Exposure", change the 30000 to 4000 and click "Set
      • If the guide probe is expected to move during the WFS exposure, then the WFS exposure will be blocked. There should be some red on the GCS GUI Guide Control window. Click
      • "Set" again to set again the desired exposure time and "unblock" the WFS.
      • Time ~ 20min to complete
    • IT 7642 - make sure NS longslit acquisitions are consistent - bypassing WCS - this should be tested during acquisition of above scripts
  6. ESM LUCI-2 Dither Test -
    • DONE ESM-DX Only
    • The scripts are in /home/lbto/Service_Observing/E_dithertest
      • E_Feb17UT_dithertest.xml (OT program file)
      • Scripts_BinoSL+L2ESM_14h10d
      • Script to use: L2_ESM_14h10d_fixed.xml
      • Time to complete ~45min (w/overheads)
  7. D-time science - LUCI-Bino QSO rest-frame H-band observations ~ low priority - This can be used as a "filler" for time between other tasks - target is above 30deg limit all night
    • Scripts are located in: /home/observer/Service_Observing/E_Poor_Weather_HbandLibrary
    • Readme file: Rothberg.H-band_Library.readme
    • Filenames: PG1444+407_L1.xml, PG1444+407_L2.xml or PG1444+407_Bino.xml and Tellurics are: Telluric_for_XXXXX
    • Finderchart is in directory
    • QSO can be observed in better conditions than the library (Observations can be executed in seeing (DIMM) up to ~1.5" and up to 3 mags of clouds)
      • Target is PG 1444+407 14:46:45 +40:35:05
      • Observations are G210 0.5" slit K-band filter - Binocular
      • If one LUCI fails - continue in monocular mode
      • Time to complete ~1hour including Telluric
  8. LUCI-Bino Non-LInearity Test - 30min (needs clear skies) - target is abvoe 30 elevation at 11:30pm
    • IC 4665 - H-band
    • Scripts are located in: /home/observer/Service_Observing/E_NonLin_Test
      • Readme file: NonLinTest.Readme
      • IC4665_Apr22_2019.gif - visibility plot
      • IC4665_NonLinTest_L1.xml (LUCI-1 Monocular), IC4665_NonLinTest_L2.xml (LUCI-2 Monocular), IC4665_NonLinTest_Bino.xml (BinoLUCI), LUCI_NonLin_Test_OnSky.xml - OT XML file
      • Coords: 17 46 18.00 +05 43 00
      • Time to complete: ~30min (w/overheads)
  9. On-Sky checkout of LUCI-1/SX/SOUL - LUCI-2/DX ESM - M92 - 1-2 hours
    • 20-min needed for each mag of AO Ref star (2 stars - 40min)
    • Need to collect on-sky data for DX AO using new KLv26
    • Test of SX SOUL - stability - including dithering, AO Ref stars (R=9, R=15)
    • Collection of ESM-SL, ESM-ESM, data of same field
    • Test of AO (SX) vs ESM (DX) for R=15 AO ref star to measure corrections as a function of distance
  10. Poor weather if AO loops can't be closed - and/or other tests can't be run - H-band Library
    • Scripts are located in: /home/observer/Service_Observing/E_Poor_Weather_HbandLibrary
    • Readme file: Rothberg.H-band_Library.readme
      • 11 science targets (late-type stars)
      • Visibility across the entire night
    • If one LUCI fails, switch to mono for the other LUCI and continue.
    • Each star+telluric takes ~30 min to complete w/overheads - if S/N is low in spectra - repeat dither pattern ~ binocular completion
  11. Low-priority filler science program - CO Velocity dispersion of coalesced mergers - observations can be done until 2am - after this point, we should do ESM/AO regardless of conditions (except where we cannot close loop - then switch to H-band Library)
    • Scripts are located in /home/observer/Service_Observing/D_Rothberg_Mergers/
    • Readme file: BR_Merger_CO_Bandhead.README
    • If one LUCI fails, switch to mono for other LUCI and continue
    • Each target takes ~50min w/Telluric for Binocular - 1hr20min w/Telluric for mono

      • Targets: RA Dec

      • UGC 5101 09 35 51 61 21 11
      • NGC 3921 11 51 06 55 04 43
      • NGC 3656 11 23 38 53 50 30

-- BarryRothberg - 29 Apr 2019
Topic attachments
I Attachment Action Size Date Who Comment
DX_20190505_RPOL.txttxt DX_20190505_RPOL.txt manage 895 K 06 May 2019 - 03:03 JohnHill John's DX IDL capture
SX_20190505_LPOL.txttxt SX_20190505_LPOL.txt manage 264 K 06 May 2019 - 03:03 JohnHill John's SX IDL capture
Topic revision: r26 - 08 May 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback